OpenStudy (anonymous):

What type of linear system is 3x − 18y = 20 and 6y − x = –12? inequivalent system inconsistent system dependent system independent system

OpenStudy (anonymous):

Hey I remember you

OpenStudy (anonymous):

I guess you got to solve for x

OpenStudy (anonymous):

what grade is this

OpenStudy (anonymous):

This is precalculas 12th grade math

OpenStudy (anonymous):

from the 17th ward

OpenStudy (anonymous):

see i cant do that

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@texaschic101

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hey

OpenStudy (misty1212):

lets solve it and see

OpenStudy (misty1212):

oops

OpenStudy (anonymous):

you was good I know what you meant

OpenStudy (misty1212):

\[3x-18y=20\\ -x+6y=-12\]

OpenStudy (misty1212):

these two lines have the same slope, but they are not the same line, they are parallel if you multiply the second equation by \(-3\) you get the two equations \[3x-18y=20\\ 3x-18y=36\]

OpenStudy (misty1212):

there is no way for \(3x-18y\) to be both \(20\) AND \(36\) so there is no solution, i.e. they are parallel

OpenStudy (anonymous):

so they are ineq

OpenStudy (anonymous):

inequivalent

OpenStudy (misty1212):

i believe the math teachers call this "inconsistent"

OpenStudy (misty1212):

i better look that up

OpenStudy (anonymous):

you right :)

OpenStudy (anonymous):

thank you so much

OpenStudy (misty1212):

nothing like google huh \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

nothing like it

OpenStudy (misty1212):

and you are welcome!